OpenStudy (anonymous):

Zen wants to fill a window box with soil. The length of the window box is 5 feet, width is 6 inches, and height is 9 inches. Each bag of soil contains 0.625 cubic foot of soil. How many bags of soil will Zen need to fill the window box completely? [1 foot = 12 inches]

OpenStudy (anonymous):

this is pretty much asking for the volume of the window box and then divide the volume of the window box by .625

OpenStudy (anonymous):

yeah but idk how to do it

OpenStudy (anonymous):

volume of box = (length) (width) (height) = 5 * 6 * 9 = 270 how many bags of 0.625 can go into 270? 270/0.625 =

OpenStudy (anonymous):

whch u get 432

OpenStudy (anonymous):

5*6*9=270 cubic feet for the box. 270/.625=432 bags

OpenStudy (anonymous):

@sithrebel15 you have to convert all the measurements into either feet or inches. Reread the problem

OpenStudy (anonymous):

60*6*9=3240 then 3240 is the volume so u divide it by .625 which is 5184

OpenStudy (anonymous):

so the answer is 5184 @Oceans

OpenStudy (anonymous):

No because they forgot to move the bags to inches too. so you go 60*6*9=3240. Then you have to convert the .625 feet to inches.

OpenStudy (anonymous):

so the answer is .625 feet

OpenStudy (anonymous):

oh my b im blind

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

yea its 432

OpenStudy (campbell_st):

its V = 5 x 0.5 x 0.75 thats it... you need all units the same.

OpenStudy (campbell_st):

the volume is V = 1.875 cubic ft

OpenStudy (campbell_st):

so the number of values is bags = volume /0.625 or bags = 1.875/0.625 so it looks like 3 bags to me